OpenStudy (anonymous):

Bradley and Ryan are having a contest to see who can collect the most bumper stickers. Bradley started with 23 bumper stickers and plans to add 4 each month. Ryan has 9 bumper stickers and plans to add 6 each month. After how many months will they have the same number of bumper stickers? A. 10 months B. 2 months C. 7 months D. 51 months

OpenStudy (anonymous):

C. 7 months

OpenStudy (anonymous):

Because: 23+(7)*4=9+(7)*6 23+28=9+42 51=51

OpenStudy (anonymous):

7 months

OpenStudy (anonymous):

i figured i was right! thanks!